OpenStudy (sharkfinsoup):

so for x^2+5 divided by x+1 is the remainder x+5? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

You should get a single number (not a variable expression) for the remainder. Try again

jimthompson5910 (jim_thompson5910):

hint: you can use the remainder theorem x+1 = x-(-1) is in the form x-k for any polynomial p(x), if we divide this by x-k, then r = p(k) is our remainder

OpenStudy (sharkfinsoup):

 that was how I did it the first time around